If Sam and odel were telling pizzas and sam sold half as many pizzas as odel and together they sold 126 pizzas. How many pizzas did sam sell? (That is the question] i tried multipying "x(x/2)=126" but the answer didnt come out right. how would you set this equation up to solve the problem ?
Answer by edjones(8007)

You had the right idea but you add the numbers instead multiplying.
Sam's pizzas + Odel's = 126
Let x=number of pizzas Odel sold.
x/2 + x = 126
x+2x=252 multiply each side by 2 to eliminate fraction.
3x=252
x=84
84/2=42 Number of pizzas Sam sold.
.
Ed